### v1

# SynParaSpeech: Automated Synthesis of Paralinguistic Datasets for Speech Generation and Understanding

Abstract
Paralinguistic sounds, like laughter and sighs, are crucial for synthesizing more realistic and engaging speech. However, existing methods typically depend on proprietary datasets, while publicly available resources often suffer from incomplete speech, inaccurate or missing timestamps, and limited real-world relevance. To address these problems, we propose an automated framework for generating large-scale paralinguistic data and apply it to construct the SynParaSpeech dataset. The dataset comprises 6 paralinguistic categories with 118.75 hours of data and precise timestamps, all derived from natural conversational speech. Our contributions lie in introducing the first automated method for constructing large-scale paralinguistic datasets and releasing the SynParaSpeech corpus, which advances speech generation through more natural paralinguistic synthesis and enhances speech understanding by improving paralinguistic event detection. The dataset and audio samples are available at https://github.com/ShawnPi233/SynParaSpeech.
Index Terms—  paralinguistic, speech synthesis, speech understanding, dataset
1 Introduction
With rapid advances in deep learning, fields such as text-to-speech (TTS) and speech-language models (SLMs) have achieved high-quality speech synthesis. Most previous methods focus on semantic content and often ignore paralinguistic sounds. Yet in natural conversation, sounds like laughter and sighs are common. Recently, more researchers have aimed to improve paralinguistic speech synthesis to enhance interactivity and realism [1, 2, 3, 4].
Although the mentioned methods can synthesize paralinguistic speech, they rely on proprietary datasets with paralinguistic annotations, which are not publicly available, limiting scalable research. Thus, open paralinguistic datasets are essential. As shown in Table 1, current open source datasets can be divided into two types. The first type includes audio-only datasets of paralinguistic events, such as AudioSet [5], ESC-50 [6], VocalSound [7], and Nonspeech7k [8]. These datasets, designed for sound event recognition, cover various non-semantic events but lack speech, text, and precise timestamps, making them unsuitable for paralinguistic speech synthesis or event localization. The second type is paralinguistic speech datasets with both speech and text, including Switchboard [9], Fisher Speech [10], MagicData-RAMC [11], NVS[12], and NVSpeech [13]. While these provide transcriptions and paralinguistic labels, they have some limitations. Switchboard and Fisher Speech have low sampling rates, Fisher Speech and MagicData-RAMC cover few event categories, and Switchboard and NVSpeech lack precise timestamps. NVS consists mainly of animations, films, and shows whose non-verbal expressions are exaggerated and over-represented, so it departs from naturally occurring conversation. Similarly, both NVS and NVSpeech construct paralinguistic datasets by training ASR models with manually annotated paralinguistic events and subsequently employing these models to label new recordings. However, such ASR-based approaches suffer from two major limitations: the imbalance of paralinguistic categories, which introduces bias into the training data, and the scarcity and high cost of paralinguistic speech data, which restrict scalability.
To address these challenges, we propose an automated approach for synthesizing large-scale paralinguistic speech data and construct the SynParaSpeech dataset. Unlike previous ASR-based expansion methods [12, 13], our approach uses paralinguistic audio events and non-performative speech recordings to automatically generate a large dataset with precise timestamp annotations. This method can be easily extended to various paralinguistic categories and languages, offering insights for building large-scale multilingual paralinguistic datasets. Our contributions are as follows:
  1. 1.
    To the best of our knowledge, we propose the first automated approach for synthesizing large-scale paralinguistic speech datasets.
  2. 2.
    We introduce SynParaSpeech, a Chinese speech dataset containing 6 paralinguistic categories, with precise timestamped transcriptions and a total duration of 118.75 hours.
  3. 3.
    We demonstrate the effectiveness of SynParaSpeech by fine-tuning CosyVoice2 and F5-TTS, achieving substantial improvements in paralinguistic speech generation.
  4. 4.
    We apply prompt tuning to models such as Qwen 2.5 Omni and Kimi Audio, enhancing the detection of paralinguistic events, and explore the impact of varying prompt context lengths.
Table 1: Paralinguistic Datasets Comparison.
Dataset Hours Clips Type Lang. SR (kHz) Timestamps Speech Available
AudioSet [5] 72.3 26,088 18 - - × ×
ESC-50 [6] 0.33 240 10 - 22.05 × ×
VocalSound [7] 20.46 21,024 6 - 16/44.1 × ×
Nonspeech7k [8] 6.75 7,014 7 - 32 × ×
Switchboard [9] 260 11,699 42 En 8 ×
Fisher Speech [10] 984 5,850 2 En 8
MagicData-RAMC [11] 180 219,325 3 Zh 16
NVS[12] 131 38,718 10 Zh/En 24
NVSpeech [13] 573.4 174,179 18 Zh - ×
SynParaSpeech (Ours) 118.75 79,986 6 Zh 24
2 Methods
Fig. 1: Overview of SynParaSpeech. (1) Labeled text with paralinguistic timestamps is synthesized. (2) Audio is synthesized and aligned with paralinguistic information. (3) The synthesized audio is verified across multiple dimensions. (4) Paralinguistic speech is generated via optimized modeling. (5) Paralinguistic speech understanding is achieved with in-context prompting.
Below, we describe how we built our dataset and the paralinguistic speech generation and understanding systems based on it. The construction of the dataset includes creating paralinguistic labeled text, synthesizing paralinguistic speech audio, and manual checks, as shown in Fig. 1.
2.1 SynParaSpeech Labeled Text Synthesis
First we designed a processing pipeline to generate accurate transcriptions aligned timestamps and paralinguistic tags. The audio was processed in two parallel steps: (1) three ASR models Whisper Large V3 [14], SenseVoice [15] and Paraformer [16] were applied to obtain sentence-level transcriptions through majority voting. (2) Voice Activity Detection (VAD) and segmented the audio into shorter clips.
Since ASR performs poorly on very short clips, we validated each VAD-determined split point by splitting the audio at these points into left and right sub-clips, transcribing each sub-clip using the same ASR models, and calculating the edit distance between these sub-clip transcriptions and the full-sentence transcription to determine how the sub-clip texts aligned within the complete sentence. For each split point, we took the transcription of the longer sub-clip, computed its Character Error Rate (CER) against the corresponding segment of the full-sentence transcription, and deemed the split accurate (yielding timestamped text clips) if the CER was below 0.1. Final alignment then leveraged these VAD outputs and ASR results, with verification via Stable Whisper [17], to generate precisely timestamped text segments.
To add paralinguistic tags, we used a mainstream large language model (LLM) Deepseek Chat V3 [18] to build a tagged text dataset. We input the aforementioned transcriptions into the LLM, prompting it to select the optimal tag from [laugh], [sigh], [gasp], [throat clearing], [pause], [tsk] without altering the original text. The LLM also found optimal insertion spots in each VAD-split clip (e.g., sentence start/clip end). Finally, we got text data with paralinguistic labels.
2.2 SynParaSpeech Audio Synthesis
The objective of the audio synthesis phase (Stage II) is to generate speech containing paralinguistic cues that conforms to the paralinguistic-annotated text produced in Stage I, by integrating established synthesis methodologies for paralinguistic audio and speech audio.
First, using the paralinguistic label from Stage I (e.g., [laugh] as in Fig. 1), we randomly pick an audio clip from the matching paralinguistic audio corpus. To keep timbre consistent between paralinguistic audio and speech, we conduct voice conversion (VC): paralinguistic audio is the source, speech audio the target. We use ASR model Whisper Large V3 [14] to encode the source audio’s semantic content, and speaker encoder CAM++ [19] to extract the target’s speaker embedding. These features go into zero-shot VC model SeedVC [20] to get paralinguistic audio with adjusted timbre.
Additionally, speech audio is sliced based on the timestamped text from Stage I. Then, the timbre-converted paralinguistic audio is inserted into the corresponding positions of the speech segments. Finally, all processed speech segments are merged in chronological order to get the final speech.
2.3 Manual Auxiliary Verification
Although the proposed process enables efficient automated synthesis of paralinguistic speech, additional manual verification was conducted to ensure alignment with human perception in terms of naturalness, audio quality, and paralinguistic categories. Professionals in speech were invited to evaluate and refine the synthesized SynParaSpeech. The evaluation considered four aspects: the naturalness of the audio, with emphasis on fluent transitions between speech and paralinguistic cues; the quality of paralinguistic expressions, including timbre consistency and accurate label matching; the overall audio quality, ensuring the absence of noise, clipping, or distortion; and the accuracy of temporal alignment between audio and text, avoiding missing, incorrect, or redundant characters. After refinement, only compliant audio clips were retained.
3 SynParaSpeech Dataset
Table 2: Statistics of the SynParaSpeech Dataset.
Category Hours Clips Avg.(s) Share
Sigh 28.22 17,706 5.74 23.76 %
Throat Clearing 25.45 18,827 4.87 21.43 %
Laugh 20.84 13,023 5.76 17.55 %
Pause 18.30 9,643 6.83 15.41 %
Tsk 14.82 11,941 4.47 12.48 %
Gasp 11.11 8,846 4.52 9.36 %
Total 118.75 79,986 5.34 100.00 %
Table 3: Paralinguistic TTS with SynParaSpeech dataset. Confidence interval for MOS scores is 95%.
Model PMOS $\uparrow$ NMOS $\uparrow$ SMOS $\uparrow$ QMOS $\uparrow$ CER(%) $\downarrow$ SECS $\uparrow$ UTMOSv2 $\uparrow$
F5-TTS (Baseline) 1.16 ± 0.01 4.08 ± 0.02 4.52 ± 0.02 3.95 ± 0.03 6.01 0.76 3.01
     + NVS SFT 1.49 ± 0.03 3.83 ± 0.03 4.03 ± 0.02 3.75 ± 0.03 12.56 0.74 3.01
     + SynParaSpeech SFT 3.10 ± 0.04 4.16 ± 0.02 4.41 ± 0.02 4.08 ± 0.02 7.26 0.74 2.83
CosyVoice2 (Baseline) 1.88 ± 0.04 4.24 ± 0.02 3.71 ± 0.03 4.00 ± 0.03 6.58 0.70 3.13
     + NVS SFT 2.35 ± 0.05 4.06 ± 0.02 3.47 ± 0.03 3.95 ± 0.03 9.50 0.69 3.02
     + SynParaSpeech SFT 3.31 ± 0.04 4.11 ± 0.02 3.74 ± 0.03 4.01 ± 0.02 11.00 0.71 2.78
     + DPO-Staged 3.40 ± 0.04 4.15 ± 0.02 3.84 ± 0.02 4.09 ± 0.02 10.91 0.70 2.87
     + DPO-Joint 3.46 ± 0.04 4.17 ± 0.02 4.03 ± 0.03 4.12 ± 0.02 11.78 0.71 2.83
As shown in Table 2, the SynParaSpeech dataset includes speech data for six paralinguistic events: sigh, throat clearing, laugh, pause, tsk, and gasp. Its statistical metrics cover total duration (Hours), number of clips (Clips), average clip duration (Avg.(s)), and proportion in the full dataset (Share) for each event.
Since the natural distribution of semantic scenarios affects paralinguistic event frequency, we did not enforce equal proportions for each category during dataset construction. Even so, statistics still show good balance across all categories: the highest proportion (sigh, 23.76%) and lowest proportion (gasp, 9.36%) differ within a reasonable range, and no category is overly dominant or extremely scarce (sample sizes range from 8,846 to 18,827). Thus, the dataset can provide balanced multi-category support for paralinguistic event analysis and modeling.
4 Experiments
4.1 Experimental Setups
To evaluate the effectiveness of SynParaSpeech in both paralinguistic speech synthesis and speech understanding tasks, we conducted experiments on paralinguistic TTS and event detection.
4.1.1 Paralinguistic TTS
In the paralinguistic TTS experiments, we employed SynParaSpeech to conduct supervised fine-tuning (SFT) on state-of-the-art open-source TTS models, namely the autoregressive CosyVoice2 [1] and the non-autoregressive F5-TTS [21]. For comparison, we also considered the NVS dataset [12], a large and diverse paralinguistic resource collected from real-world media and automatically annotated. To ensure comparability, we selected four common paralinguistic labels—[laugh], [sigh], [gasp], and [throat clearing]—that are supported by both SynParaSpeech and NVS. We kept 2% of the data randomly as a validation set, the rest for training. For evaluation, the LLM DeepSeek V3 [18] was used to rewrite seed-tts-eval texts (with paralinguistic labels inserted), making 100 inference texts per category.
We compared the CosyVoice2 and F5-TTS checkpoints with their fine-tuned derivatives trained on SynParaSpeech and on NVS. To overcome the issue that SFT only focuses on the model’s overall output loss and lacks focus on paralinguistic synthesis, we applied direct preference optimization (DPO) [22] to CosyVoice2 in two configurations: DPO-staged (DPO after SFT) and DPO-joint (optimizing both objectives together). Leveraging the clean origin of SynParaSpeech, each utterance naturally provides paired positives (with paralinguistic events) and negatives (without). During DPO training, the original utterance serves as the rejected sample $\mu_{\mathrm{o}}$, while its SynParaSpeech counterpart is the chosen sample $\mu_{\mathrm{p}}$. The preference learning objective is thus defined as
$\Delta(\mu_{\mathrm{p}},\mu_{\mathrm{o}};y)$ $=\log\frac{\pi_{\mathrm{ref}}(\mu_{\mathrm{p}}\mid y)}{\pi_{\theta}(\mu_{\mathrm{p}}\mid y)}-\log\frac{\pi_{\mathrm{ref}}(\mu_{\mathrm{o}}\mid y)}{\pi_{\theta}(\mu_{\mathrm{o}}\mid y)},$ (1)
$\mathcal{L}_{\text{dpo}}$ $=-\mathbb{E}_{y}\left[\log\sigma\left(\beta\cdot\Delta(\mu_{\mathrm{p}},\mu_{\mathrm{o}};y)\right)\right],$
where $\pi_{\theta}$ is the target model, $\pi_{\mathrm{ref}}$ the reference model, $\beta=0.01$ the temperature, and $\sigma(\cdot)$ the sigmoid function.
For training, CosyVoice2’s language model is optimized with Adam at a learning rate of $1\times 10^{-5}$ for 50 epochs, with early stopping (patience 10), a constant learning rate, and a 2500-step linear warm-up. Training stability is maintained via gradient clipping (threshold 5), gradient accumulation over 2 steps, and dynamic batching with a maximum of 2000 frames. Following the NVS setup, F5-TTS is trained for 400 epochs with a learning rate of $1\times 10^{-4}$, a cosine annealing scheduler with 1000 warm-up updates, and a frame-based batch size of 30,000 per GPU. To initialize new paralinguistic tokens, we use embeddings from the RoBERTa-based [23] text encoder of CLAP [24], aligning dimensions via interpolation.
Model performance was assessed using both objective and subjective metrics. Objective measures include character error rate (CER) for intelligibility, UTMOSv2 [25] for speech quality, and SECS for timbre similarity. Subjective evaluation employed 5-point ratings: SMOS for speaker similarity, NMOS for naturalness, QMOS for overall audio quality, and PMOS for the paralinguistic quality. For MOS ratings, 21 volunteers participated in a double-blind evaluation.
4.1.2 Paralinguistic Event Detection
Table 4: Paralinguistic event detection results.
Model Context Acc. $\uparrow$ F1 Score $\uparrow$ CER (%) $\downarrow$
Kimi Audio - 0.320 0.294 17.79
1 0.314 0.312 11.30
3 0.354 0.336 10.61
5 0.382 0.340 11.11
7 0.371 0.331 11.01
Qwen 2.5 Omni - 0.215 0.189 23.52
1 0.337 0.357 21.18
3 0.460 0.447 20.60
5 0.473 0.471 19.48
7 0.423 0.362 20.07
In the paralinguistic event detection experiments, we applied SynParaSpeech-based prompt tuning to multimodal large language models (MLLMs) and analyzed the effect of context size. Following MMSU [26], Kimi Audio [27] was chosen for its strong paralinguistic reasoning, and Qwen 2.5 Omni [28] for its superior paralinguistic perception. Both models were evaluated with and without SynParaSpeech prompt tuning. For each paralinguistic category, 100 audio–text pairs were used for testing, while 1, 3, 5, and 7 pairs were sampled as context prompts. Performance was measured using accuracy, macro F1 score, and CER.
4.2 Experimental Results and Analysis
4.2.1 Paralinguistic TTS Experimental Results
The experimental results of paralinguistic TTS are shown in Table 3. Comparing F5-TTS across the two datasets, the model fine-tuned on SynParaSpeech shows improvements in paralinguistic synthesis, naturalness, and overall quality. The gain is most evident in paralinguistic synthesis, where PMOS increases significantly compared with the model fine-tuned on NVS. SynParaSpeech also provides advantages in naturalness, speaker similarity, overall quality, and articulation. Similarly, CosyVoice2 fine-tuned on SynParaSpeech achieves a large improvement in paralinguistic synthesis, a moderate gain in speaker similarity, and overall advantages over the NVS-fine-tuned counterpart in paralinguistic quality, naturalness, speaker similarity, and overall quality. These results confirm that SynParaSpeech can enhance paralinguistic synthesis while preserving naturalness, speaker similarity, and overall speech quality.
Furthermore, F5-TTS uses non-autoregressive flow matching, it achieves high naturalness and speaker similarity. In contrast, CosyVoice2, with its autoregressive multi-stage design, better models paralinguistic features and produces higher audio quality. Building on this, we investigated how DPO enables CosyVoice2 to better capture paralinguistic details and further enhance its paralinguistic capability. The last three rows show that DPO boosts paralinguistic quality, naturalness, and speaker similarity. For both DPO types, our joint SFT training beats staged training, increases all subjective scores, and achieves the best PMOS and QMOS.
We also observed that models trained on paralinguistic datasets show slight declines in objective metrics CER, SECS, and UTMOSv2, as these metrics are designed for standard speech. For example, laughter may be transcribed as “ha ha”, raising CER. Nevertheless, models trained on SynParaSpeech still maintain high NMOS and QMOS while achieving clear PMOS gains.
4.2.2 Paralinguistic Event Detection Experimental Results
Table 4 highlights the effect of SynParaSpeech prompt tuning and context quantity on paralinguistic event detection. First, the results confirm the utility of SynParaSpeech: for both Kimi Audio and Qwen 2.5 Omni, introducing SynParaSpeech prompts substantially improves performance compared with the no-context baseline, especially in terms of accuracy and macro F1 score. This verifies that the SynParaSpeech dataset provides effective supervision signals for paralinguistic perception and reasoning.
Second, the number of contextual examples exerts a critical influence. Performance generally rises with the increase in contextual quantity, and peak performance gains are observed under the 5-shot setting for both models. However, excessive context (e.g., 7-shot) fails to bring further improvements and may even lead to performance degradation. This indicates a trade-off between providing informative prompts and avoiding model overload caused by redundant input.
Overall, these findings demonstrate that SynParaSpeech enhances the model’s capability in paralinguistic event detection, while an optimal quantity of contextual examples is essential to fully leverage its advantages.
5 Conclusion
This paper proposes an automated approach for large-scale paralinguistic dataset synthesis and introduces the SynParaSpeech dataset. SynParaSpeech, featuring multiple paralinguistic categories with fine-grained annotations and a substantial scale, serves as an effective resource to facilitate both paralinguistic speech synthesis (TTS) and paralinguistic event detection tasks. Experimental results demonstrate that integrating SynParaSpeech enhances the generation quality of TTS models and boosts the performance of paralinguistic event detection models.